OpenStudy (anonymous):

Prove the Pythagorean theorem using diagrams and the following facts: the area of a square with side s is s^2 and the area of a triangle with base b and height h is 1/2bh

OpenStudy (zzr0ck3r):

The area of the ouside square is (a+b)^2 another way to find the area is to find the are of each triangle add them up and then add the area of the inner box to that so (a+b)^2=4(1/2)ba+c^2 a^2+2ab+b^2=2ab+c^2 a^2+b^2=c^2